RC meeting and Run Plan, Friday, April 4th
From SBS wiki
Contents
RC Meeting Information
Current RC: Bogdan Wojtsekhowski
RC Daily Meetings are at 1:30 (pm) via Zoom and in person in the counting house meeting room (2nd floor).
- Zoom connection link: https://jlab-org.zoomgov.com/j/1616220064
- Meeting ID: 161 622 0064
- Pass code: 314159
Useful Links
Accelerator
- Accelerator delivered beam to Hall C, tuning continue.
- At 11am today CEBAF is OFF due to cooling water leak which used for the power supplies, https://logbooks.jlab.org/entry/4348041
- Hall A start likely Tuesday, 4/8, defined by target cooldown schedule.
Completed Work
- Beamline vacuum leak - fixed by D.Meekins
- CDet: DAQ made good progress
- GEM: DAQ taking data with HCAL
Work planned/ongoing in/around the Hall
- Target preparation
- DAQ development
- Detector testing
Shifts
- Shifts canceled through Tuesday OWL (4/8)
Status of Subsystem
- GEp documentation and permits,wiki pageECal,cosmic at 1500 V https://logbooks.jlab.org/entry/4347614,plan for heating testHCal,signal vs. 1 GeV, cosmic calibrationGEMS,DAQ testsCDET,cosmic hit distribution https://logbooks.jlab.org/entry/4347567DAQAnalysis Software,online monitoring, 50k replay
Current Issues
- Beam line: https://logbooks.jlab.org/entry/4347555
- Trigger latency in each DAQ esp. CDET
- Experts to update wiki for shift takers (DAQ/GEM/CDET/HCAL/ECAL/50k replay)
- Detailed run plan (PI meeting on Monday)
- NEW time for RC meeting: 1:30 pm for Tu, Th, Fri and 1pm for Mo and We.
Beam Restoration Procedure
- Use new "cross target" (1x1 mm "+" shape) for beam restoration. https://logbooks.jlab.org/entry/4346117
Once we have beam
Beam requirements
- start off at 3 pass beam
- intrinsic beam size: 150–500 µm
- raster size: 2 mm X 2 mm; using carbon hole target with 5mm hole to checkout raster size
- For later
- production beam current : 50 µA
- run plan forthcoming
SBS Magnet
- First the first 3 pass, SBS magnet will be at 41.5% of maximum
- corrector magnet setting for each pass passed to Yves Roblin